RSS Feed
News
Nov
10
[RESOLVED] Connectivity Issues for Virgin Media customers
Posted by Phil Parry on 10 November 2016 10:23 AM

We are seeing problems with Virgin Media customers who are having problems reaching our services as of around 9am this morning.

This is NOT a problem with our network or services  - the issue lies wholly within Virgin Media network.

* Update * Our network department have made changes to route around the problem.

* RESOLVED * Our networks department have reported that the issue that caused issues for Virgin Media customers is resolved.  

The issue was caused as a result of problems within the LINX peering exchange that many major ISPs use to exchange network traffic.


Read more »



Nov
10
[RESOLVED] Issues affecting SQL 10 server
Posted by Jonathan Hughes on 10 November 2016 08:58 AM

Engineers are working to resolve an error with the SQL 10 server.

Please accept our apologies for the outage and any difficulties caused by this.

update: 10:20 The issue is now resolved and sql service running normally. 


Read more »



Oct
24
[ MAINTENANCE ] Shared Hosting Linux Plesk and cPanel server restarts.
Posted by Anthony Hogbin on 24 October 2016 11:42 AM


As you may be aware there has been a lot of press regarding the zero day vulnerability Dirty CoW this weekend.

We have been working on this since Friday - and have moved from our temporary measures to stable solutions today.

This morning a long term stable patch has been applied to our Linux shared Plesk and cPanel nodes, and a restart followed this.

As a result you may have experienced a loss of service as these machines came back up.

If you continue to have service issues - please do not assume this is related. These works have been completed now - so any residual issues are likely unrelated - so get them raised with support@hostinguk.net .

If you have any questions regarding this Copy on Write vulnerability - do not hesitate to get in touch with us.

This maintenance window is now closed.


Read more »



Oct
7
[ Maintenance Window ] Upgrades to Cloud / VM platform.
Posted by Anthony Hogbin on 07 October 2016 10:49 AM

Engineering work will be occurring during the following window -

Day: Sunday 9th October.

Time: 2100hrs BST (UTC+1).

During a time window of: 3 Hours.

Expected downtime:Less than 1 hour.

Activity:

HUK staff and platform vendors will be running through a series of platform upgrades to our visualization platform. During this period the control panel will be unavailable. As each host node is restarted and bought up with new images - there will an interruption to service as instances restart. A full hardware and software health check will be carried out on completion of works. Services will be interrupted by a reboot.

We will update this as the works progress.


Read more »



Sep
23
[ LOSS OF PROTECTION ] 0000-0500hrs Tuesday 27th September.
Posted by Anthony Hogbin on 23 September 2016 12:35 PM

Day: Morning of Tuesday 27th September;

Time: Midnight to 5am Europe/London;

Reason: Soft and firmware maintenance along redundant transit route;

Effect: Loss of protection to core fibre and resulting transit to DC8 / St Asaph.

Description:

DC8 is supplied by redundant geodiverse fibre feeds.

During the early hours of Tuesday 27th - our transit providers will be carrying out work on a section of our reundant leg.

They have allowed a 3 hour window of hard down while software and firmware updates are applied, services restarted and tested.

We do not envisage any effect to services what so ever - however we will lose the N+1 protection for this period.

As such no services or transit will be effected directly.

Completion:

We will update on safe completion.

[UPDATE 1109]

This test had been carried out - but has not completed well.

THIS HAS NOT EFFECTED ANY OF OUR HARDWARE OR SERVICES.

Post testing a modular UPS/PDU dropped the load causing a core switch failure.

We will remain in loss of protection state until this has been remedied.

[UPDATE 1300]

This engineering has completed successfully, and is to be considered closed.


Read more »